OpenStudy (anonymous):

(1,-1) ; parallel to y=2/5x-3 ! help

OpenStudy (anonymous):

If two lines are parallel, they have the same slope. That means the slope of the line will also be 2/5. We also have a point on the line: (1,-1). Using these two facts you can get the equation of the line by taking y=mx+b. m is just the slope 2/5 x and y come from the point (1,-1) We can use this to solve for b. -1=(2/5)*1+b b=-7/5 so y= 2/5x - 7/5